# Evaluating diversion and treatment policies for opioid use disorder

### v3

@@ 2.5 Calibration and Validation of Model Outputs
For model calibration and validation, the simulation was run 600 times with different random seeds to construct 95% joint prediction intervalss for the main simulation outputs of interest. We also estimate monetary costs for each of these states to approximate the cost savings, not including implementation costs, of a given policy. We use the common random numbers method to reduce variation in the simulation model when using each distribution in Table 1.
The model's structure in Figure 1 was validated by consulting expert opinion via incorporating feedback from both the subject matter experts in policing and medicine. The policing subject matter expert provided high, low, and best guess estimates for the number of individuals in the Dane County jail system and $p_{A}$ in Table 1. The medical subject matter expert provided high, low, and best guess estimates for the number of individuals currently in OUD treatment and the hospital due to opioid-related hospital encounters. Calibration adjustments were made to input parameters that relied on expert opinion (Law,, 2005). Appendix C reports additional model outputs that were reviewed and validated by the subject matter experts in policing and addiction treatment.
@@
Figure 2 shows the model validation, where the resulting calibrated model output PIs are compared to the actual value of the corresponding target Dane County data. As shown in Figure 2, the warm-up period was determined when the yearly outputs were similar to the 2013 to 2017 outcomes observed in the data sources. Only the pre-2018 target data points were used to estimate calibration error, since an arrest diversion and overdose diversion program were implemented in Dane County at the end of 2017 (Zgierska et al.,, 2021; Wisconsin Voices For Recovery,, 2021). Therefore, we use 17 calibration targets to calibrate the simulation. The targets are the number of OUD treatments from 2014 to 2017, the number of opioid-related deaths from 2013 to 2017, the number of opioid-related arrests from 2015 to 2017, and the number of opioid-related hospital encounters from 2013 to 2017, which are summarized as “target data” in Figure 2. We describe how the target data is collected in Appendix A.5. The goodness-of-fit measures used are the number of target points within the simulation model 95% PIs. The search algorithm we used was multi-approached. First, we adjusted the subject matter expert estimates by $\pm$ 5% of the corresponding inputs that lay outside the 95% PIs. Secondly, we used an alternative calculation of our parameter estimates. Additional details regarding our final parameter estimation are described in Appendix A. As shown in Figure 2, the model has an average calibration error of 5.88% among the 17 target Dane County data points, where only the number of Dane County hospital encounters in 2014 fell outside of the 95% PI.
@@ 2.6 Validation of Cost Estimates
We verified the accuracy of the simulation model’s societal cost estimates by comparing the total simulated costs for 2017 with the estimated societal costs for Dane County. Table 3 summarizes the Dane County 2017 number of events and estimated total societal costs, as well as the simulated 2017 95% PIs of total events and total simulated societal costs. We present a societal cost average and range for the active use state, representing the lowest and highest proportions observed in other states (e.g., OUD treatment, hospital encounter) multiplied by the estimated number of cases of opioid misuse in 2017 for Wisconsin (i.e., 205,000) (SAMHSA, SAMHSA). This, therefore, calculates a mean and range for the total estimated 2017 number of active users at year-end and the associated societal costs for Dane County. To determine the total simulated societal costs for a specific scenario, we multiply the USD per event costs for 2017 by the yearly 95% PI simulated event estimates, resulting in a 95% PI total yearly cost estimate. As shown in Table 3, the 2017 estimated costs are within each of the 95% PIs of each simulated cost. The simulated active use at year-end and the hospital encounter-related costs are near the lower end of the 95% PIs, while opioid-related death and opioid-related arrest costs are at the high end. Overall, the total societal costs of all states combined are on the low end of the estimated costs in 2017. This means a given policy scenario’s actual cost savings are likely higher than the simulated values.
@@
Active use at year-end 17,630 (12,505, 28,085) (14,072, 15,909) \$601.3 (\$426.5, \$957.9) (\$479.9, \$542.6)
Opioid-related death 88 (50, 101) \$1,016.3 (\$577.4, \$1166.4)
Opioid-related arrest 545 (456, 617) \$30.4 (\$25.4, \$34.4)
Start OUD treatment 2026 (1692, 2100) \$3.7 (\$3.1, \$3.8)
Hospital encounter 1718 (1,720, 2,054) \$34.5 (\$34.4, \$41.1)
Total cost: \$1,686 (\$1,511, \$2,043) (\$1,144, \$1,765)
@@ 3 Results and Discussion
Table 4 reports several cumulative outputs through 2032. The year 2032 is ten years from the potential implementation of the CM policy and 16 years from the potential implementation of the AD and OD policies. Table 4 reports 95% CI standard errors and mean difference p-values for the cumulative number of opioid-related deaths, non-diverted opioid-related arrests, opioid hospital encounters, OUD treatment episodes, and individuals in the active use state at year end. Additionally, Table 4 reports each scenario’s simulated total costs, cost difference from the base model, and cost mean difference p-values. A two-sided paired t-test of the mean difference from the base model (i.e., (0, 22, 0)) is used to report statistically significant differences. P-values in boldface with a () are below a significance level of 0.05, and those in boldface with an (∗∗) are below a level of 0.001. All cumulative outputs in Table 4 have a p-value below 0.05.
Table 4: Simulated mean $\pm$ standard error (p-value) of cumulative total events, cost, and savings for scenarios after 10 years: 2023 through 2032
@@
Base Model 0, 22, 0 2169.81 $\pm$ 2.82 14386.08 $\pm$ 14.62 51860.08 $\pm$ 50.71 51211.01 $\pm$ 51.62 404280.65 $\pm$ 378.59 42846.08 $\pm$ 46.1 0.0
Arrest Diversion 20, 22, 0 2166.53$\pm$ 2.78 (0.004) 12372.91$\pm$ 13.23∗∗ ($<$0.001) 51817.81$\pm$ 50.66∗∗ ($<$0.001) 53167.61$\pm$ 52.74∗∗ ($<$0.001) 403959.39$\pm$ 378.48∗∗ ($<$0.001) 42672.72$\pm$ 45.36∗∗ ($<$0.001) -173.36
40, 22, 0 2166.17$\pm$ 2.76 (0.002) 10366.14$\pm$ 12.08∗∗ ($<$0.001) 51773.35$\pm$ 50.82∗∗ ($<$0.001) 55121.67$\pm$ 54.06∗∗ ($<$0.001) 403612.71$\pm$ 378.39∗∗ ($<$0.001) 42540.95$\pm$ 45.2∗∗ ($<$0.001) -305.13
60, 22, 0 2163.99$\pm$ 2.78∗∗ ($<$0.001) 8369.32$\pm$ 10.74∗∗ ($<$0.001) 51731.28$\pm$ 50.65∗∗ ($<$0.001) 57066.61$\pm$ 55.26∗∗ ($<$0.001) 403279.26$\pm$ 378.58∗∗ ($<$0.001) 42383.19$\pm$ 45.21∗∗ ($<$0.001) -462.89
80, 22, 0 2162.66$\pm$ 2.81∗∗ ($<$0.001) 6370.61$\pm$ 9.51∗∗ ($<$0.001) 51690.86$\pm$ 50.68∗∗ ($<$0.001) 59015.2$\pm$ 56.42∗∗ ($<$0.001) 402962.42$\pm$ 378.85∗∗ ($<$0.001) 42238.94$\pm$ 45.88∗∗ ($<$0.001) -607.14
100, 22, 0 2160.24$\pm$ 2.8∗∗ ($<$0.001) 4378.23$\pm$ 8.2∗∗ ($<$0.001) 51651.35$\pm$ 50.85∗∗ ($<$0.001) 60951.65$\pm$ 57.69∗∗ ($<$0.001) 402655.02$\pm$ 378.88∗∗ ($<$0.001) 42081.3$\pm$ 45.38∗∗ ($<$0.001) -764.78
Overdose Diversion 0, 30, 0 2165.74$\pm$ 2.74 (0.041) 14363.32$\pm$ 14.54∗∗ ($<$0.001) 51782.95$\pm$ 50.7∗∗ ($<$0.001) 53854.18$\pm$ 53.11∗∗ ($<$0.001) 403665.47$\pm$ 378.3∗∗ ($<$0.001) 42777.73$\pm$ 45.53 (0.007) -68.35
0, 45, 0 2155.71$\pm$ 2.72∗∗ ($<$0.001) 14318.87$\pm$ 14.56∗∗ ($<$0.001) 51629.37$\pm$ 50.62∗∗ ($<$0.001) 58934.47$\pm$ 56.49∗∗ ($<$0.001) 402387.02$\pm$ 378.95∗∗ ($<$0.001) 42609.98$\pm$ 44.83∗∗ ($<$0.001) -236.1
0, 60, 0 2150.95$\pm$ 2.72∗∗ ($<$0.001) 14275.62$\pm$ 14.56∗∗ ($<$0.001) 51478.21$\pm$ 50.6∗∗ ($<$0.001) 63977.57$\pm$ 59.77∗∗ ($<$0.001) 401161.88$\pm$ 377.09∗∗ ($<$0.001) 42518.37$\pm$ 44.98∗∗ ($<$0.001) -327.71
0, 75, 0 2145.68$\pm$ 2.63∗∗ ($<$0.001) 14235.08$\pm$ 14.51∗∗ ($<$0.001) 51325.94$\pm$ 50.54∗∗ ($<$0.001) 68977.86$\pm$ 62.76∗∗ ($<$0.001) 399944.09$\pm$ 377.54∗∗ ($<$0.001) 42407.25$\pm$ 43.86∗∗ ($<$0.001) -438.83
0, 90, 0 2139.51$\pm$ 2.68∗∗ ($<$0.001) 14189.87$\pm$ 14.43∗∗ ($<$0.001) 51178.12$\pm$ 50.4∗∗ ($<$0.001) 73939.78$\pm$ 65.88∗∗ ($<$0.001) 398734.28$\pm$ 375.97∗∗ ($<$0.001) 42283.29$\pm$ 44.19∗∗ ($<$0.001) -562.79
Case Management 0, 22, 20 2165.57$\pm$ 2.81∗∗ ($<$0.001) 14364.62$\pm$ 14.6∗∗ ($<$0.001) 51777.48$\pm$ 50.69∗∗ ($<$0.001) 54119.38$\pm$ 52.52∗∗ ($<$0.001) 403696.85$\pm$ 378.18∗∗ ($<$0.001) 42772.84$\pm$ 46.09∗∗ ($<$0.001) -73.24
0, 22, 40 2160.69$\pm$ 2.75∗∗ ($<$0.001) 14342.44$\pm$ 14.58∗∗ ($<$0.001) 51696.14$\pm$ 50.69∗∗ ($<$0.001) 57001.89$\pm$ 53.77∗∗ ($<$0.001) 403066.53$\pm$ 378.17∗∗ ($<$0.001) 42688.89$\pm$ 45.37∗∗ ($<$0.001) -157.2
0, 22, 60 2158.12$\pm$ 2.75∗∗ ($<$0.001) 14319.99$\pm$ 14.57∗∗ ($<$0.001) 51617.45$\pm$ 50.82∗∗ ($<$0.001) 59865.76$\pm$ 55.25∗∗ ($<$0.001) 402464.88$\pm$ 378.76∗∗ ($<$0.001) 42638.84$\pm$ 45.21∗∗ ($<$0.001) -207.25
0, 22, 80 2155.04$\pm$ 2.76∗∗ ($<$0.001) 14299.08$\pm$ 14.51∗∗ ($<$0.001) 51537.63$\pm$ 50.75∗∗ ($<$0.001) 62706.37$\pm$ 56.36∗∗ ($<$0.001) 401866.46$\pm$ 377.98∗∗ ($<$0.001) 42583.68$\pm$ 45.44∗∗ ($<$0.001) -262.4
0, 22, 100 2152.68$\pm$ 2.77∗∗ ($<$0.001) 14278.44$\pm$ 14.52∗∗ ($<$0.001) 51457.55$\pm$ 50.71∗∗ ($<$0.001) 65525.88$\pm$ 57.47∗∗ ($<$0.001) 401275.95$\pm$ 377.96∗∗ ($<$0.001) 42529.48$\pm$ 45.47∗∗ ($<$0.001) -316.6
Policy Mix 20, 40, 20 2153.37$\pm$ 2.74∗∗ ($<$0.001) 12315.28$\pm$ 13.23∗∗ ($<$0.001) 51564.26$\pm$ 50.54∗∗ ($<$0.001) 61829.57$\pm$ 57.81∗∗ ($<$0.001) 401950.06$\pm$ 377.67∗∗ ($<$0.001) 42457.07$\pm$ 45.41∗∗ ($<$0.001) -389.01
40, 60, 40 2141.07$\pm$ 2.64∗∗ ($<$0.001) 10278.06$\pm$ 12.03∗∗ ($<$0.001) 51263.56$\pm$ 50.38∗∗ ($<$0.001) 72538.52$\pm$ 63.55∗∗ ($<$0.001) 399553.18$\pm$ 376.21∗∗ ($<$0.001) 42104.82$\pm$ 43.9∗∗ ($<$0.001) -741.26
60, 80, 60 2130.45$\pm$ 2.75∗∗ ($<$0.001) 8280.5$\pm$ 10.79∗∗ ($<$0.001) 50975.51$\pm$ 50.11∗∗ ($<$0.001) 82601.65$\pm$ 69.26∗∗ ($<$0.001) 397266.47$\pm$ 374.78∗∗ ($<$0.001) 41787.17$\pm$ 45.42∗∗ ($<$0.001) -1,058.92
*Statistically significant difference from the Base Model, with p-value $<0.05$, using 2 -sided paired t-test with a level of 0.05
**Statistically significant difference from the Base Model, with p-value $<0.001$, using 2-sided paired t-test with a level of 0.05
@@
At the 0.001 level, Table 4 indicates that a policy that has at least 60% AD, 45% OD, or 20% CD statistically significantly decreases the number of non-diverted opioid-related arrests, opioid-related deaths, active opioid use, and hospital encounters while statistically significantly increasing the amount of OUD treatment episodes. At the 0.05 level, all policies show statistically significant cumulative reductions to the main outputs. This suggests that arrest diversion, overdose diversion, and case management can lead to cumulative reductions in opioid-related arrests, opioid-related deaths, hospitalizations, and opioid use with any amount of policy implementation. The differences in these outputs are greater when mixing multiple policies and at higher levels of policy implementation. For example, with the CM-focused policies of (0, 22, 20) with 20% CM compared to (0, 22, 60) with 60% CM, we see a decrease in the expected average number of opioid-related arrests from 14,365 to 14,320, respectively. If we combine multiple policies, we see an even larger reduction in opioid-related arrests to 12,315.28 for the policy-mix (20, 40, 20). Therefore, when jurisdictions plan to implement a treatment policy, they should develop a plan to implement multiple policies over time and scale each policy within their communities, such as offering a treatment policy to more people.

### v4

@@ 2.5 Calibration and Validation of Model Outputs
For model calibration and validation, the simulation was run 600 times with different random seeds to construct 95% joint prediction intervals for the main simulation outputs of interest. We also estimate monetary costs for each of these states to approximate the cost savings, not including implementation costs, of a given policy. We use the common random numbers method to reduce variation in the simulation model when using each distribution in Table 1.
The model's structure in Figure 1 was validated by consulting expert opinion via incorporating feedback from both the subject matter experts in policing and medicine. The policing subject matter expert provided high, low, and best guess estimates for the number of individuals in the Dane County jail system and $p_{A}$ in Table 1. The medical subject matter expert provided high, low, and best guess estimates for the number of individuals currently in OUD treatment and the hospital due to opioid-related hospital encounters. Appendix C reports additional model outputs that were reviewed and validated by the subject matter experts in policing and addiction treatment.
@@
Figure 2 shows the model validation, where the resulting calibrated model output PIs are compared to the actual value of the corresponding target Dane County data. As shown in Figure 2, the warm-up period was determined when the yearly outputs were similar to the 2013 to 2017 outcomes observed in the data sources. Only the pre-2018 target data points were used to estimate calibration error, since an arrest diversion and overdose diversion program were implemented in Dane County at the end of 2017 (Zgierska et al.,, 2021; Wisconsin Voices For Recovery,, 2021). Therefore, we use 17 calibration targets to calibrate the simulation. The targets are the number of OUD treatments from 2014 to 2017, the number of opioid-related deaths from 2013 to 2017, the number of opioid-related arrests from 2015 to 2017, and the number of opioid-related hospital encounters from 2013 to 2017, which are summarized as “target data” in Figure 2. We describe how the target data is collected in Appendix A.5. The goodness-of-fit measures used are the number of target points within the simulation model 95% PIs. The search algorithm we used was multi-approached. First, we adjusted the subject matter expert estimates by $\pm$ 5% of the corresponding inputs that lay outside the 95% PIs. Parameters calibrated via this first method are denoted in Table 1 with “C” as they were calibrated with an initial estimation that relied on expert opinion (Law,, 2005). Secondly, we used an alternative calculation of our parameter estimates. Parameters calibrated via this second method are denoted in Table 1 with “DM”, since they have multiple direct estimations that are considered. Additional details regarding our final parameter estimation are described in Appendix A. As shown in Figure 2, the model has an average calibration error of 11.76% among the 17 target Dane County data points, where only the number of Dane County hospital encounters in 2014 and 2017 fell outside of the 95% PI.
@@ 2.6 Validation of Cost Estimates
We verified the accuracy of the simulation model’s societal cost estimates by comparing the total simulated costs for 2017 with the estimated societal costs for Dane County. Table 3 summarizes the Dane County 2017 number of events and estimated total societal costs, as well as the simulated 2017 95% PIs of total events and total simulated societal costs. We present a societal cost average and range for the active use state, representing the lowest and highest proportions observed in other states (e.g., OUD treatment, hospital encounter) multiplied by the estimated number of cases of opioid misuse in 2017 for Wisconsin (i.e., 205,000) (SAMHSA, SAMHSA). This, therefore, calculates a mean and range for the total estimated 2017 number of active users at year-end and the associated societal costs for Dane County. To determine the total simulated societal costs for a specific scenario, we multiply the USD per event costs for 2017 by the yearly 95% PI simulated event estimates, resulting in a 95% PI total yearly cost estimate. As shown in Table 3, the total 2017 estimated costs are within the 95% PIs of the simulated cost. The simulated active use at year-end and the hospital encounter-related costs are near the lower end of the 95% PIs, while opioid-related death and opioid-related arrest costs are at the high end. Overall, the total simulated societal costs of all states combined are on the low end of the estimated costs in 2017. This means a given policy scenario’s actual cost savings are likely higher than the simulated cost values.
@@
Active use at year-end 17,630 (12,505, 28,085) (14,277, 16,218) \$601.3 (\$426.5, \$957.9) (\$486.9, \$553.1)
Opioid-related death 88 (48, 104) \$1016.3 (\$554.3, \$1201)
Opioid-related arrest 545 (452, 639) \$30.4 (\$25.2, \$35.6)
Start OUD treatment 2026 (1751, 2149) \$16.7 (\$14.4, \$17.7)
Hospital encounter 1718 (1749, 2102) \$20.7 (\$21.1, \$25.3)
Total cost: \$1686 (\$1511, \$2042) (\$1,115, \$1,809)
@@ 3 Results and Discussion
3.1 10-Year Cumulative Policy Events: 2023 through 2032
Table 4 reports several 10-year cumulative outputs from the start of 2023 through the end of 2032.
Table 4: Simulated mean $\pm$ standard error (p-value) of 10-year cumulative total events, cost, and savings for scenarios: 2023 through 2032
@@
Base Model 0, 22, 0 1038.02 $\pm$ 1.42 6706.75 $\pm$ 4.99 23573.9 $\pm$ 14.4 24589.71 $\pm$ 15.64 192270.75 $\pm$ 112.0 21,478.0 $\pm$ 20.57 0.0
Arrest Diversion 20, 22, 0 1036.17 $\pm$ 1.44(0.063) 5358.7$\pm$ 4.28∗∗ ($<$0.001) 23553.95$\pm$ 14.37∗∗ ($<$0.001) 25909.08$\pm$ 16.75∗∗ ($<$0.001) 192105.33$\pm$ 112.68∗∗ ($<$0.001) 21,376.0$\pm$ 20.73∗∗ ($<$0.001) -102.61∗∗
40, 22, 0 1036.52 $\pm$ 1.43(0.135) 4014.55$\pm$ 3.37∗∗ ($<$0.001) 23531.22$\pm$ 14.4∗∗ ($<$0.001) 27216.7$\pm$ 17.3∗∗ ($<$0.001) 191890.41$\pm$ 112.5∗∗ ($<$0.001) 21,300.0$\pm$ 20.97∗∗ ($<$0.001) -178.4∗∗
60, 22, 0 1033.43$\pm$ 1.44∗∗ ($<$0.001) 2672.26$\pm$ 2.56∗∗ ($<$0.001) 23504.16$\pm$ 14.19∗∗ ($<$0.001) 28525.93$\pm$ 17.79∗∗ ($<$0.001) 191696.84$\pm$ 111.35∗∗ ($<$0.001) 21,186.0$\pm$ 20.96∗∗ ($<$0.001) -292.19∗∗
80, 22, 0 1033.71$\pm$ 1.43∗∗ ($<$0.001) 1335.86$\pm$ 1.61∗∗ ($<$0.001) 23484.69$\pm$ 14.33∗∗ ($<$0.001) 29828.8$\pm$ 18.72∗∗ ($<$0.001) 191504.22$\pm$ 111.73∗∗ ($<$0.001) 21,110.0$\pm$ 20.71∗∗ ($<$0.001) -368.34∗∗
100, 22, 0 1032.11$\pm$ 1.47∗∗ ($<$0.001) 0.0$\pm$ 0.0∗∗ ($<$0.001) 23459.58$\pm$ 14.31∗∗ ($<$0.001) 31135.28$\pm$ 19.47∗∗ ($<$0.001) 191313.94$\pm$ 112.04∗∗ ($<$0.001) 21,013.0$\pm$ 21.16∗∗ ($<$0.001) -465.0∗∗
Overdose Diversion 0, 30, 0 1038.34 $\pm$ 1.4(0.849) 6693.43$\pm$ 4.97∗∗ ($<$0.001) 23531.48$\pm$ 14.31∗∗ ($<$0.001) 26359.38$\pm$ 16.66∗∗ ($<$0.001) 191900.29$\pm$ 111.31∗∗ ($<$0.001) 21,490.0 $\pm$ 20.45(0.61) 11.44
0, 45, 0 1027.26$\pm$ 1.38∗∗ ($<$0.001) 6670.23$\pm$ 5.05∗∗ ($<$0.001) 23453.72$\pm$ 14.29∗∗ ($<$0.001) 29778.07$\pm$ 18.68∗∗ ($<$0.001) 191221.15$\pm$ 112.24∗∗ ($<$0.001) 21,344.0$\pm$ 20.15∗∗ ($<$0.001) -134.75∗∗
0, 60, 0 1026.97$\pm$ 1.44∗∗ ($<$0.001) 6645.81$\pm$ 4.98∗∗ ($<$0.001) 23372.19$\pm$ 14.23∗∗ ($<$0.001) 33166.83$\pm$ 20.71∗∗ ($<$0.001) 190496.64$\pm$ 111.65∗∗ ($<$0.001) 21,339.0$\pm$ 20.94∗∗ ($<$0.001) -139.51∗∗
0, 75, 0 1023.01$\pm$ 1.35∗∗ ($<$0.001) 6620.47$\pm$ 4.96∗∗ ($<$0.001) 23289.89$\pm$ 14.14∗∗ ($<$0.001) 36529.44$\pm$ 22.39∗∗ ($<$0.001) 189767.75$\pm$ 111.32∗∗ ($<$0.001) 21,281.0$\pm$ 19.85∗∗ ($<$0.001) -197.41∗∗
0, 90, 0 1020.43$\pm$ 1.42∗∗ ($<$0.001) 6598.3$\pm$ 4.97∗∗ ($<$0.001) 23210.16$\pm$ 14.15∗∗ ($<$0.001) 39876.39$\pm$ 24.42∗∗ ($<$0.001) 189060.21$\pm$ 111.06∗∗ ($<$0.001) 21,250.0$\pm$ 20.59∗∗ ($<$0.001) -228.2∗∗
Case Management 0, 22, 20 1034.81$\pm$ 1.44∗∗ ($<$0.001) 6686.57$\pm$ 5.0∗∗ ($<$0.001) 23494.89$\pm$ 14.27∗∗ ($<$0.001) 27532.79$\pm$ 16.96∗∗ ($<$0.001) 191668.86$\pm$ 111.61∗∗ ($<$0.001) 21,440.0$\pm$ 20.79 (0.003) -38.65∗∗
0, 22, 40 1029.91$\pm$ 1.46∗∗ ($<$0.001) 6662.63$\pm$ 5.01∗∗ ($<$0.001) 23412.31$\pm$ 14.37∗∗ ($<$0.001) 30449.56$\pm$ 18.28∗∗ ($<$0.001) 191041.48$\pm$ 112.21∗∗ ($<$0.001) 21,377.0$\pm$ 21.22∗∗ ($<$0.001) -101.72∗∗
0, 22, 60 1027.84$\pm$ 1.44∗∗ ($<$0.001) 6641.69$\pm$ 5.02∗∗ ($<$0.001) 23330.99$\pm$ 14.21∗∗ ($<$0.001) 33341.59$\pm$ 19.49∗∗ ($<$0.001) 190453.61$\pm$ 111.12∗∗ ($<$0.001) 21,353.0$\pm$ 20.55∗∗ ($<$0.001) -125.56∗∗
0, 22, 80 1023.12$\pm$ 1.47∗∗ ($<$0.001) 6620.54$\pm$ 4.94∗∗ ($<$0.001) 23251.73$\pm$ 14.18∗∗ ($<$0.001) 36208.12$\pm$ 20.8∗∗ ($<$0.001) 189845.6$\pm$ 110.68∗∗ ($<$0.001) 21,300.0$\pm$ 21.39∗∗ ($<$0.001) -178.81∗∗
0, 22, 100 1020.25$\pm$ 1.4∗∗ ($<$0.001) 6598.45$\pm$ 4.92∗∗ ($<$0.001) 23173.76$\pm$ 14.14∗∗ ($<$0.001) 39057.07$\pm$ 22.01∗∗ ($<$0.001) 189257.63$\pm$ 111.38∗∗ ($<$0.001) 21,259.0$\pm$ 20.53∗∗ ($<$0.001) -219.85∗∗
Policy Mix 20, 40, 20 1026.72$\pm$ 1.37∗∗ ($<$0.001) 5319.9$\pm$ 4.27∗∗ ($<$0.001) 23381.8$\pm$ 14.34∗∗ ($<$0.001) 32609.93$\pm$ 19.92∗∗ ($<$0.001) 190681.75$\pm$ 111.51∗∗ ($<$0.001) 21,267.0$\pm$ 20.56∗∗ ($<$0.001) -211.32∗∗
40, 60, 40 1020.28$\pm$ 1.44∗∗ ($<$0.001) 3956.64$\pm$ 3.39∗∗ ($<$0.001) 23191.32$\pm$ 14.07∗∗ ($<$0.001) 40510.23$\pm$ 23.65∗∗ ($<$0.001) 189075.59$\pm$ 109.85∗∗ ($<$0.001) 21,105.0$\pm$ 20.86∗∗ ($<$0.001) -373.84∗∗
60, 80, 60 1009.8$\pm$ 1.44∗∗ ($<$0.001) 2618.82$\pm$ 2.5∗∗ ($<$0.001) 23023.47$\pm$ 14.08∗∗ ($<$0.001) 47828.35$\pm$ 27.81∗∗ ($<$0.001) 187662.98$\pm$ 110.06∗∗ ($<$0.001) 20,894.0$\pm$ 21.18∗∗ ($<$0.001) -584.39∗∗
@@
The year 2032 is ten years from the potential implementation of the CM policy and 16 years from the potential implementation of the AD and OD policies. Table 4 reports 95% CI standard errors and mean difference p-values for the 10-year cumulative number of opioid-related deaths, non-diverted opioid-related arrests, opioid hospital encounters, OUD treatment episodes, and individuals in the active use state at year end. Additionally, Table 4 reports each scenario’s simulated total costs, cost difference from the base model, and cost mean difference p-values. A two-sided paired t-test of the mean difference from the base model (i.e., (0, 22, 0)) is used to report statistically significant differences. P-values in boldface with a () are below a significance level of 0.05, and those in boldface with an (∗∗) are below a level of 0.001.
@@
At the 0.001 level, Table 4 indicates that a policy that has at least 60% AD, 45% OD, or 20% CM statistically significantly decreases the number of non-diverted opioid-related arrests, opioid-related deaths, active opioid use, and hospital encounters while statistically significantly increasing the amount of OUD treatment episodes. All policies show statistically significant cumulative reductions to non-diverted opioid-related arrests, active opioid use, and hospital encounters, suggesting that arrest diversion, overdose diversion, and case management can lead to cumulative reductions in opioid-related arrests, hospitalizations, and opioid use with any amount of policy implementation. The differences between all policy outputs, including opioid-related deaths, from the base model are greater when mixing multiple policies and at higher levels of policy implementation. For example, with the CM-focused policies of (0, 22, 20) with 20% CM compared to (0, 22, 60) with 60% CM, we see a decrease in the expected average number of opioid-related arrests from 6,687 to 6,642, respectively. If we combine multiple policies, we see an even larger reduction in non-diverted opioid-related arrests to 5,320 for the policy-mix (20, 40, 20). Therefore, when jurisdictions plan to implement a treatment policy, they should develop a plan to implement multiple policies over time and scale each policy within their communities, such as offering a treatment policy to more people.
3.2 Societal Costs